### v1

# Collapses in quantum-classical probabilistically checkable proofs and the quantum polynomial hierarchy

Abstract
We investigate structural properties of quantum proof systems by establishing several collapse results that uncover simplifications in their complexity landscape. By extending the classical results such as the Karp–Lipton theorem to quantum polynomial hierarchy with quantum proofs and establishing uniqueness preservation for quantum-classical probabilistically checkable proof systems, we address how different constraints affect the computational power of these proof systems.
Our main contributions are:
  • We prove that restricting quantum-classical PCP systems to uniqueness does not reduce computational power: $\mathsf{UniqueQCPCP}=\mathsf{QCPCP}$ under $\mathsf{BQ}$-operator and randomized reductions. This equivalence demonstrates robustness, as is shown by $\mathsf{UniqueQCMA=QCMA}$ under randomized reductions result, showing that uniqueness constraints do not fundamentally limit the capabilities even in these $\mathsf{PCP}$-type proof systems.
  • We establish a non-uniform quantum analogue of the Karp–Lipton theorem, concluding a conditional “inclusive collapse” for the quantum polynomial hierarchy ($\mathsf{QPH}$): if $\mathsf{QMA}\subseteq\mathsf{BQP}/\mathsf{qpoly}$, then $\mathsf{QPH}\subseteq\mathsf{Q\Sigma}_{2}/\mathsf{qpoly}$. This extends the classical collapse theorem to quantum complexity with quantum advice, providing evidence against efficient quantum advice for $\mathsf{QMA}$-complete problems.
  • We introduce a consistent variant of the quantum polynomial hierarchy ($\mathsf{CQPH}$) that enforces consistency constraints across interaction rounds while maintaining product‑state proofs. We prove its unconditional collapse to the second level, i.e., $\mathsf{CQPH}=\mathsf{CQ\Sigma}_{2}$ showing that only consistency constraints restrict the computational power of higher levels in the hierarchy. We emphasize that this is in contrast to prior work on quantum-entangled polynomial hierarchy ($\mathsf{QEPH}$), which credits $\mathsf{QEPH}$ collapse to entanglement.
These results contribute to our understanding of structural boundaries in quantum complexity theory and the interplay between different constraint types in quantum proof systems.
1 Introduction
Computational complexity theory seeks to understand the fundamental limits and capabilities of computational models under various resource constraints. A particularly rich and fruitful direction within this field involves studying how classical computational machines can be augmented with different forms of auxiliary information to enhance their problem-solving capabilities. This auxiliary information may take various forms: deterministic advice strings that depend only on the input length, non-deterministic certificates or witnesses that can be verified efficiently, or interactive proofs where a computationally unbounded prover attempts to convince a computationally unbounded verifier of the truth of a statement through a sequence of message exchanges.
Over the past few decades, complexity theorists have made remarkable progress in characterizing the computational power of such augmented systems. The overall theme of this research program can be condensed in the following fundamental question:
“Given a particular verifier, how powerful can the system become when provided with additional information encoded as proofs or advice, and how do restrictions on these forms of information affect the verifier’s power?”
This question has led to several landmark results that have shaped our understanding of the complexity of computational models. Perhaps one of the most surprising results is the complete characterization of interactive proof systems. Goldwasser, Micali, and Rackoff initially introduced the concept of interactive proofs [GMR85], demonstrating that interaction between a prover and verifier could yield computational advantages beyond traditional proof systems. This work was extended by Babai [Bab85] with the introduction of Arthur-Merlin games, providing an alternative perspective on interactive computation. The pinnacle of this line of research came with Shamir’s groundbreaking result [Sha92], which established that the class of problems admitting interactive proofs with three or more rounds of interaction is exactly as powerful as $\mathsf{PSPACE}$. Remarkably, this result also showed that increasing the number of interaction rounds beyond three does not enhance the computational power of the system—a surprising collapse that demonstrates the efficiency of interactive protocols. The story takes a dramatic turn when we consider quantum entanglement in multi-prover interactive proofs. While classical multi-prover interactive proofs are bounded by $\mathsf{NEXP}$, the quantum variant exhibits fundamentally different behavior. In a recent breakthrough, Ji, Natarajan, Vidick, Wright, and Yuen [JNV+21] proved that $\mathsf{MIP^{*}}=\mathsf{RE}$, showing that multi-prover interactive proofs with quantum entangled provers can decide any recursively enumerable language. This result is particularly striking because it demonstrates that quantum entanglement allows the proof system to transcend the polynomial hierarchy entirely, reaching the computability-theoretic boundary of decidable problems. The contrast between the classical collapse to $\mathsf{PSPACE}$ and the quantum expansion to $\mathsf{RE}$ illustrates the profound impact that quantum resources can have on interactive proof systems.
Another cornerstone result that illustrates the delicate relationship between computational classes and auxiliary information is the Karp–Lipton theorem [KL80]. This theorem establishes a conditional collapse of the polynomial hierarchy under the assumption that $\mathsf{NP}\subseteq\mathsf{P}/\mathsf{poly}$—that is, if every problem in $\mathsf{NP}$ can be solved in polynomial time with polynomial-sized advice strings. The theorem demonstrates that such an inclusion would imply $\mathsf{PH}=\mathsf{\Sigma}_{2}$, collapsing the entire polynomial hierarchy to its second level. Since most complexity theorists believe the polynomial hierarchy is infinite (and thus does not collapse), the Karp–Lipton theorem provides strong evidence against the existence of polynomial-sized circuits for $\mathsf{NP}$-complete problems.
The 1990s witnessed another revolutionary development with the proof of the PCP (Probabilistically Checkable Proof) theorem by Arora, Lund, Motwani, Sudan, and Szegedy [ALM+98], building on earlier work by Babai, Fortnow, and Lund [BFL91] and Feige, Goldwasser, Lovász, Safra, and Szegedy [FGL+96]. This celebrated result asserts that every problem in $\mathsf{NP}$ admits a probabilistically checkable proof system with remarkable efficiency properties: specifically, any $\mathsf{NP}$ statement has a proof that can be verified by reading only a constant number of randomly chosen bits from the proof string, using only $\mathcal{O}(\log(n))$ random bits in total. The PCP theorem not only provided profound insights into the structure of $\mathsf{NP}$ but also established fundamental connections between computational complexity and approximation algorithms, showing that many optimization problems cannot be approximated within certain factors unless $\mathsf{P}=\mathsf{NP}$.
With the advent of quantum computing, researchers have naturally sought to understand quantum analogues of these classical results. The quantum setting introduces fundamentally new phenomena—such as superposition, entanglement, and measurement—that can potentially provide computational advantages while also introducing novel technical challenges. Quantum complexity classes such as $\mathsf{BQP}$ (bounded-error quantum polynomial time), $\mathsf{QMA}$ (quantum Merlin-Arthur), and their variants have been extensively studied, revealing a rich landscape of quantum computational complexity.
In the quantum setting, proof systems take on new dimensions of complexity. Quantum proofs can exist in superposition states, verifiers can perform quantum computations, and the interaction between classical and quantum information becomes a central consideration. The study of quantum probabilistically checkable proofs ($\mathsf{QPCP}$), quantum interactive proofs, and quantum advice has revealed both similarities to and fundamental differences from their classical counterparts.
1.1 Related works
Quantum probabilistically checkable proofs and unique variants.
Our first result builds upon the Valiant-Vazirani theorem [VV86], which shows a randomized reduction from $\mathsf{NP}$ to its unique variant $\mathsf{UniqueNP}$. A quantum-classical analog was developed by Aharonov, Ben-Or, Brandão and Sattath [ABOBS22], who proved a reduction from $\mathsf{QCMA}$ to $\mathsf{UniqueQCMA}$ based on the same hashing technique. Our work generalizes this approach to establish a reduction from $\mathsf{QCPCP}$ to $\mathsf{UniqueQCPCP}$ under $\mathsf{BQ}$-operator and randomized reductions.
The complexity class $\mathsf{QPCP}$ was first formally defined by Aharonov, Arad, Landau and Vazirani [AALV10], though the proof-checking formulation has received limited attention since then. Recent work by Buhrman, Helsen, and Weggemans [BHW24] and Buhrman, Le Gall, and Weggemans [BGW24] has established foundational results for both $\mathsf{QPCP}$ and $\mathsf{QCPCP}$. The latter work introduces the $\mathsf{BQ}$-operator to formally capture the “quantumness” inherent in $\mathsf{QCPCP}$.
In related developments, Irani, Natarajan, Nirkhe, Rao and Yuen [INN+22] demonstrated that a single query to a classical $\mathsf{PP}$ oracle suffices to prepare a $\mathsf{QMA}$ witness to inverse-polynomial accuracy. They also proved that no polynomial-time reduction from $\mathsf{QMA}$-search to $\mathsf{QMA}$-decision exists relative to a suitable quantum oracle. Their “state synthesis” framework resolved Aaronson’s question [Aar16] regarding sublinear query complexity for arbitrary $n$-qubit states.
Quantum polynomial hierarchy and related complexity classes.
The quantum polynomial hierarchy was first defined by Yamakami [Yam03], where the verifier uses alternating quantum proofs. The standard description used in this paper differs from the original definition in that we consider quantum circuits rather than quantum Turing machines. Gharibian and Kempe [GK12] subsequently studied $\mathsf{cq\text{-}\Sigma}_{2}$ (more commonly denoted $\mathsf{QC\Sigma}_{2}$) with universally quantified proofs.
Lockhart and Guillén [LG17] introduced a hierarchy similar to $\mathsf{QCPH}$ but using existential and universal operators. While this hierarchy does not fully capture $\mathsf{QCMA}$, its structural properties are significantly easier to analyze than those of $\mathsf{QCPH}$, which we do not discuss in detail here.
Significant progress on the $\mathsf{QCPH}$ hierarchy was made by Gharibian, Santha, Sikora, Sundaram and Yirka [GSS+22] and Agarwal, Gharibian, Koppula and Rudolph [AGKR24]. [GSS+22] formally defined $\mathsf{QPH}$ and $\mathsf{QCPH}$ as used in this paper, and proved weaker variants of the Karp-Lipton theorem and Toda’s theorem specifically for $\mathsf{QCPH}$. As a corollary, they showed that proving a Karp-Lipton theorem for $\mathsf{QPH}$ would imply $\mathsf{QMA}(2)\subseteq\mathsf{PSPACE}$, which appears beyond current techniques. We emphasize that we do not provide a standard Karp-Lipton statement for $\mathsf{QPH}$; rather, we prove a conditional inclusion collapse under the second level of $\mathsf{QPH}$ with polynomial-length quantum advice. They also established $\mathsf{QMA}(2)\subseteq\mathsf{Q\Sigma}_{3}\subseteq\mathsf{NEXP}$, strengthening the upper bound on $\mathsf{QMA}(2)$.
The collapse of $\mathsf{QCPH}$ was proven independently by Falor, Ge and Natarajan [FGN23] and Agarwal, Gharibian, Koppula and Rudolph [AGKR24]. The latter work upper bounds $\mathsf{QCPH}$ by showing $\mathsf{QCPH}\subseteq\mathsf{pureQPH}\subseteq\mathsf{EXP^{PP}}$. Grewal and Yirka [GY24] demonstrated $\mathsf{QCPH}\subseteq\mathsf{QPH}$ at the expense of a constant factor blowup in the hierarchy level.
Agarwal, Gharibian, Koppula and Rudolph [AGKR24] proved a Karp-Lipton theorem for $\mathsf{QCPH}$ and error reduction for $\mathsf{pureQPH}$, while we establish a non-uniform version of Karp-Lipton for $\mathsf{QPH}$. Grewal and Yirka [GY24] showed an unconditional collapse of $\mathsf{QEPH}$ to its second level, proving equality with $\mathsf{QRG}(1)$. They attributed this collapse to “cheating” enabled by entanglement between consecutive rounds. However, we refine this insight by demonstrating that the consistency property between proof rounds is solely responsible for the collapse, as our consistent-$\mathsf{QPH}$ exhibits the same collapse to the second level as $\mathsf{QEPH}$. They also established $\mathsf{QCPH}=\mathsf{DistributionQCPH}$, where proofs in the latter are distributions over strings.
1.2 Our contributions
In this work, we investigate quantum analogues of two important variants of proof systems, contributing new insights to our understanding of quantum complexity theory. Our results span two main areas: quantum probabilistically checkable proofs and the quantum polynomial hierarchy.
Quantum-classical probabilistically checkable proofs.
In Section 2, we examine the relationship between general quantum-classical PCP systems ($\mathsf{QCPCP}$) and their unique variant ($\mathsf{UniqueQCPCP}$), where the proof system is required to have a unique accepting quantum proof state. We demonstrate that restricting quantum-classical PCP systems to the unique variant does not reduce their computational power under $\mathsf{BQ}$-operator and randomized reductions. This establishes a collapse of general $\mathsf{QCPCP}$ systems to their uniqueness-restricted counterpart, providing new structural insights into quantum proof systems.
Quantum polynomial hierarchy.
In Section 3, we establish two significant results concerning the quantum polynomial hierarchy ($\mathsf{QPH}$). First, we prove a non-uniform quantum analogue of the Karp–Lipton theorem: if $\mathsf{QMA}\subseteq\mathsf{BQP}/\mathsf{qpoly}$ (i.e., if every problem in quantum Merlin-Arthur can be solved by a bounded-error quantum polynomial-time machine with quantum polynomial advice), then the quantum polynomial hierarchy collapses to its second level with quantum advice: $\mathsf{QPH}\subseteq\mathsf{Q\Sigma_{2}}/\mathsf{qpoly}$. This result parallels the classical case while accounting for the unique features of quantum computation and quantum advice.
Second, we introduce and analyze a novel variant of the quantum polynomial hierarchy inspired by the entangled-prover setting studied in recent work on quantum entangled proofs [GY24]. This variant, which we term the consistent quantum polynomial hierarchy, enforces consistency conditions across different rounds of interaction while maintaining the constraint that proof states remain in product form (i.e., unentangled). We prove an unconditional collapse of this consistent version to its second level, showing that $\mathsf{CQ\Sigma}_{k}=\mathsf{CQ\Sigma}_{2}$ for all $k\geq 2$. Further, we also show similar results for their corresponding pure versions (refer Definition 3.13). This demonstrates that inter-round consistency constraints significantly restrict the computational power of such hierarchical quantum proof systems, providing new insights into the role of consistency in quantum interactive protocols.
2 Equivalence of unique and standard quantum-classical PCPs
In this section, we establish our first collapse result by proving that
[…]
2.1 Definitions and preliminaries
2.1.1 Complexity classes
We formally define the unique variants we introduce, namely $\mathsf{UniqueQCPCP}$, $\mathsf{UniqueMTP}$, and $\mathsf{UniqueSAT}$.
Definition 2.1 (Unique quantum-classical probabilistically checkable proof, $\mathsf{UniqueQCPCP}$).
Let $c,s$ be completeness and soundness parameters, and let
$p,q:\mathbb{N}\to\mathbb{N},\qquad c,s:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}.$
A promise problem $\mathcal{A}=(A_{\mathrm{YES}},A_{\mathrm{NO}})$ is in $(p(n),q(n),c,s)$-$\mathsf{UniqueQCPCP}$ if there exists a uniformly generated family of quantum circuits $\{V_{n}:n\in\mathbb{N}\}$ of polynomial size such that each $V_{n}$ receives an input $x\in\{0,1\}^{n}$ and a proof string $y\in\{0,1\}^{p(n)}$, from which at most $q(n)$ bits can be queried. The verifier accepts by measuring a designated output qubit and observing outcome $1$, and satisfies the following:
  1. (i)
    (Completeness and Uniqueness) If $x\in A_{\mathrm{YES}}$, then there exists a unique $y\in\{0,1\}^{p(n)}$ such that
    $\Pr[V_{n}(x,y)\textnormal{ accepts}]\geq c,$ (2.1)
    and for all $y^{\prime}\neq y$,
    $\Pr[V_{n}(x,y^{\prime})\textnormal{ accepts}]\leq s.$ (2.2)
  2. (ii)
    (Soundness) If $x\in A_{\mathrm{NO}}$, then for all $y\in\{0,1\}^{p(n)}$,
    $\Pr[V_{n}(x,y)\textnormal{ accepts}]\leq s.$ (2.3)
We write $\mathsf{UniqueQCPCP}[q]$ if $p(n)=\textnormal{poly}(n)$, $c=2/3$, and $s=1/3$.
Definition 2.2 (Unique multilinear threshold problem, $\mathsf{UniqueMTP}$).
Let $P:\{0,1\}^{N}\to\mathbb{R}$ be a degree-$d$ multilinear polynomial with real coefficients $\{\beta_{S}\}_{S\subseteq[N],|S|\leq d}$. Suppose the image of $P$ lies in a finite set $D\subset[0,1]$ of evenly spaced values, such that $\log_{2}|D|\leq\mathrm{poly}(N)$. Denote by $\delta_{D}$ the maximum element of $D$. Given a polynomial-size classical description of $P$ (i.e., the coefficients $\{\beta_{S}\}$), the set $D$, and a threshold $a\in D$, decide between the following two cases:
  1. (i)
    There exists a unique $y\in\{0,1\}^{N}$ such that $P(y)\geq a$ and $P(y^{\prime})<a$ for all $y^{\prime}\neq y$.
  2. (ii)
    For all $y\in\{0,1\}^{N}$, it holds that $P(y)<a$.
Definition 2.3 (Unique propositional satisfiability problem, $\mathsf{UniqueSAT}$).
Let $\phi$ be a Boolean formula over $n$ variables. Define the set of satisfying assignments as
$\mathsf{SAT}(\phi)=\{x\in\{0,1\}^{n}:\phi(x)=1\}.$ (2.4)
Then the language
$\mathsf{UniqueSAT}=\{\phi:|\mathsf{SAT}(\phi)|=1\}$ (2.5)
consists of Boolean formulas that admit exactly one satisfying assignment.
2.1.2 Reduction techniques
In this section, we present the formal definition of randomized reductions, as well as two formulations of the $\mathsf{BQ}$-operator as defined in [BGW24].
Definition 2.4 (Randomized polynomial-time reduction).
Let $\mathcal{A}$ and $\mathcal{B}$ be languages. A randomized polynomial-time reduction from $\mathcal{A}$ to $\mathcal{B}$ is a randomized algorithm $\mathcal{R}$ such that, for every input $x$:
  1. (i)
    If $x\in\mathcal{A}$, then $\Pr[\mathcal{R}(x)\in B]\geq 1/2$.
  2. (ii)
    If $x\notin A$, then $\Pr[\mathcal{R}(x)\in B]=0$.
The algorithm $\mathcal{R}$ must run in time polynomial in $|x|$.
Lemma 2.5 (Valiant–Vazirani [VV86]).
There exists a randomized polynomial-time algorithm
:{CNF formulas over $\mathcal{R}:\{\textnormal{CNF formulas over }n\textnormal{ variables}\}\rightarrow\{\textnormal{CNF formulas}\}$ (2.6)
such that, for every Boolean formula $\phi$ over $n$ variables, the following holds:
  1. (i)
    (Completeness) If $\phi$ is satisfiable (i.e., $|\mathsf{SAT}(\phi)|\geq 1$), then
    $\Pr\left[\mathcal{R}(\phi)\textnormal{ has exactly one satisfying assignment}\right]\;\geq\;\frac{1}{2n}.$ (2.7)
  2. (ii)
    (Soundness) If $\phi$ is unsatisfiable, then $\mathcal{R}(\phi)$ is unsatisfiable with probability $1$.
Moreover, $\mathcal{R}$ runs in time polynomial in $|\phi|$.
Definition 2.6 ($\mathsf{BQ}$-operator [BGW24, Definition 6]).
Let $\mathcal{C}$ be a class of promise problems. Define $\mathsf{BQ}\cdot\mathcal{C}$ as the class of all promise problems $\mathcal{A}$ for which there exists a promise problem $\mathcal{B}\in\mathcal{C}$ such that $\mathcal{A}$ reduces to $\mathcal{B}$ via a polynomial-time quantum reduction with success probability at least $2/3$:
$\mathsf{BQ}\cdot\mathcal{C}:=\{\mathcal{A}:\mathcal{A}\leq_{q}\mathcal{B}\textnormal{ for some }\mathcal{B}\in\mathcal{C}\}.$ (2.8)
Definition 2.7 (Alternative formulation of $\mathsf{BQ}$-operator [BGW24, Definition 6a]).
Let $\mathcal{C}$ be a class of promise problems. Then $\mathsf{BQ}\cdot\mathcal{C}$ consists of all promise problems $\mathcal{A}=(A_{\mathrm{YES}},A_{\mathrm{NO}})$ such that there exists a promise problem $\mathcal{B}=(B_{\mathrm{YES}},B_{\mathrm{NO}})\in\mathcal{C}$ and a polynomial-time quantum algorithm $\mathcal{Q}$ with the following properties: for all $x$ with $|x|=n$,
  1. (i)
    (Completeness) If $x\in A_{\mathrm{YES}}$, then
    $\Pr_{z\sim\mathcal{Q}(x)}\left[(x,z)\in B_{\mathrm{YES}}\right]\geq 2/3.$ (2.9)
  2. (ii)
    (Soundness) If $x\in A_{\mathrm{NO}}$, then
    $\Pr_{z\sim\mathcal{Q}(x)}\left[(x,z)\in B_{\mathrm{NO}}\right]\geq 2/3.$ (2.10)
We note that when the base class $\mathcal{C}$ contains at least the class $\mathsf{P}$, we adopt Definition 2.7 as the default interpretation of the $\mathsf{BQ}$-operator. Otherwise, Definition 2.6 is used. In this work, Definition 2.7 will suffice for all purposes.
2.2 Reduction chain
In [BGW24], a $\mathsf{BQ}$ reduction from $\mathsf{QCPCP}$ to $\mathsf{MTP}$ is presented in detail. In this section, we first show that $\mathsf{MTP}$ can be efficiently and deterministically reduced to $\mathsf{SAT}$. The reduction from $\mathsf{SAT}$ to $\mathsf{UniqueSAT}$ then follows directly from [VV86] via an randomized reduction. Subsequently, we establish a reduction from $\mathsf{USAT}$ to $\mathsf{UniqueMTP}$, and finally a $\mathsf{BQ}$ reduction from $\mathsf{UniqueMTP}$ to $\mathsf{UniqueQCPCP}$. The overall proof structure is summarized in Figure 1.
Figure 1: Reduction structure used to prove the collapse result $\mathsf{UniqueQCPCP}=\mathsf{QCPCP}$. The left side illustrates the original reduction chain from $\mathsf{QCPCP}$, as established in [BGW24]. The right side depicts our refined reduction path to the unique variant, culminating in $\mathsf{UniqueQCPCP}$ (Definition 2.1). Key components include deterministic and randomized reductions, as well as the application of the $\mathsf{BQ}$-operator.
Lemma 2.8 (Deterministic polynomial-time reduction from $\mathsf{MTP}$ to $\mathsf{SAT}$).
The $\mathsf{MTP}$ can be reduced to $\mathsf{SAT}$ deterministically and in polynomial time.
Proof.
Let $P(y)$ denote the solution polynomial of the given $\mathsf{MTP}$ instance
$P(y)=\sum_{\begin{subarray}{c}S\subseteq[p(n)],~{}|S|\leq 2q\end{subarray}}\beta_{S}\prod_{i\in S}y_{i},$ (2.11)
where $\beta_{S}\in\mathbb{R}$ is represented using $\ell$-bit fixed-point encoding.
We assume the standard promise structure:
  1. (i)
    (Correctness and Uniqueness) There exists at least a $y\in\{0,1\}^{n}$ such that $P(y)\geq a$.
  2. (ii)
    (Soundness) For all $y\in\{0,1\}^{n}$, $P(y)<a$.
Transforming $P(y)$ into a Boolean formula.
For each monomial $\prod_{i\in S}y_{i}$, define a clause:
$\phi_{S}:=\bigwedge_{i\in S}y_{i}$ (2.12)
This is a conjunction over a subset of input variables. Each such clause $\phi_{S}$ can be encoded in CNF form via the Tseytin transformation.
Each $\beta_{S}$ satisfies the bound ([BGW24, Lemma 3]):
$|\beta_{S}|\leq(1+(1+2^{2q})^{2q-1})\|P\|_{\infty},\quad\text{with }\|P\|_{\infty}\leq 1.$ (2.13)
Thus, $\beta_{S}$ can be expressed as $v_{S}/2^{\ell}$ with $v_{S}\in\mathbb{Z}$ represented using $k$-bit integers, where:
$k=\log_{2}(1+(1+2^{2q})^{2q-1})+\ell+1.$ (2.14)
Define the scaled polynomial:
$P^{\prime}(y):=2^{\ell}\cdot P(y)=\sum_{\begin{subarray}{c}S\subseteq[p(n)],~{}|S|\leq 2q\end{subarray}}v_{S}\cdot\phi_{S},$ (2.15)
where $v_{S}$ is a $k$-bit integer. Write $v_{S}$ in binary:
$v_{S}=\sum\limits_{i=1}^{k^{\prime}}2^{i}b_{S,i},$ (2.16)
with $b_{S,i}\in\{0,1\}$ and $k^{\prime}=\log_{2}(1+(1+2^{2q})^{2q-1})$.
As $\phi_{S}\in\{0,1\}$, it acts like a Boolean variable. Thus:
$P^{\prime}(y)=\sum_{\begin{subarray}{c}S\subseteq[p(n)],~{}|S|\leq 2q\end{subarray}}\sum_{i}2^{i}(b_{S,i}\land\phi_{S})=\sum_{S}\left(\sum_{i}2^{i}b_{S,i}\right)\cdot\phi_{S}.$ (2.17)
Adder network construction.
Each $v_{S}$ is represented in $k^{\prime}$ bits. To compute $P^{\prime}(y)$, we sum all such $k^{\prime}$-bit numbers. This can be done using standard half-adder and full-adder networks, which can be encoded into CNF form using the Tseytin transformation. The total circuit has size polynomial in $|S|\cdot k^{\prime}$ and logarithmic depth $\mathcal{O}(\log(|S|\cdot k^{\prime}))$.
Threshold comparison circuit.
We wish to enforce:
$\exists y\in\{0,1\}^{n}:P^{\prime}(y)\geq 2^{\ell}a,$ (2.18)
or in the soundness case,
$\forall y\in\{0,1\}^{n},\;P^{\prime}(y)<2^{\ell}a.$ (2.19)
These threshold comparisons can be implemented by comparator circuits (in the class CC) and converted to CNF via Tseytin transformations.
Putting it all together.
All parts of $P^{\prime}(y)$—the monomials, the binary representations, the adder and comparator circuits—can be encoded into a Boolean formula in CNF with auxiliary variables $b_{S,i}$ and original variables $y_{i}$. This yields a deterministic and efficient reduction from $\mathsf{MTP}$ to $\mathsf{SAT}$. ∎
Lemma 2.9 (Deterministic polynomial-time reduction from $\mathsf{UniqueSAT}$ to $\mathsf{UniqueMTP}$).
There exists a deterministic polynomial-time reduction from $\mathsf{UniqueSAT}$ to $\mathsf{UniqueMTP}$.
Proof.
Given a CNF formula $\phi$ with clauses $C_{1},C_{2},\ldots,C_{m}$ over variables $x_{1},x_{2},\ldots,x_{n}$, we construct a multilinear polynomial as follows:
$P(x_{1},\ldots,x_{n})=\sum_{j=1}^{m}\prod_{\ell\in C_{j}}\tilde{x}_{\ell},$ (2.20)
where for each literal $\ell$ in clause $C_{j}$:
$\tilde{x}_{\ell}=\begin{cases}x_{i}&\text{if }\ell=x_{i},\\ 1-x_{i}&\text{if }\ell=\neg x_{i}.\end{cases}$ (2.21)
We set the threshold to $a=m$.
  1. (i)
    (Completeness) If assignment $x$ satisfies $\phi$, then $x$ satisfies every clause $C_{j}$. For each clause $C_{j}$, at least one literal evaluates to true, making the corresponding product $\prod_{\ell\in C_{j}}\tilde{x}_{\ell}=1$. Therefore, $P(x)=m\geq a$.
  2. (ii)
    (Soundness) If assignment $x$ does not satisfy $\phi$, then $x$ violates at least one clause $C_{j}$. For any violated clause, all literals evaluate to false, making the corresponding product $\prod_{\ell\in C_{j}}\tilde{x}_{\ell}=0$. Thus, $P(x)\leq m-1<a$.
  3. (iii)
    (Uniqueness preservation) Since the reduction maps each satisfying assignment $x$ of $\phi$ to the same assignment $x$ for the polynomial $P$, uniqueness is preserved by the identity mapping.
The reduction runs in polynomial time as constructing $P$ requires examining each clause and literal once. ∎
Lemma 2.10 (BQ reduction from $\mathsf{UniqueMTP}$ to $\mathsf{UniqueQCPCP}$).
Let $\Psi=(P(y),\,a,\,\delta_{D})$ be an instance of the $\mathsf{UniqueMTP}$ over Boolean variables $y\in\{0,1\}^{m}$, where
$P(y)=\sum_{\begin{subarray}{c}S\subseteq[m],\,|S|\leq d\end{subarray}}\beta_{S}\prod_{i\in S}y_{i},$ (2.22)
and suppose that $\Psi$ satisfies the following promise:
$\begin{cases}\exists!\,y^{*}\in\{0,1\}^{m}\textnormal{ such that }P(y^{*})\geq a&\textnormal{(YES instance)},\\ \forall\,y\in\{0,1\}^{m},\;P(y)<a&\textnormal{(NO instance)},\end{cases}$ (2.23)
with a gap parameter $\delta_{D}>0$. Then there exists a polynomial-time deterministic quantum algorithm that outputs the description of a verifier $V_{\Psi}\in\mathsf{UniqueQCPCP}[p(m)=m,\,q=d,\,c,\,s]$, where
$c=\frac{a+\delta_{D}}{\sum_{j=1}^{N}|\beta_{S_{j}}|}\quad\text{and}\quad s=\frac{a-\delta_{D}}{\sum_{j=1}^{N}|\beta_{S_{j}}|},$
such that:
  • (i)
    If $\Psi$ is a YES instance with unique witness $y^{*}$, then $V_{\Psi}$ accepts exactly on the proof $y^{*}$ and rejects all other $y$.
  • (ii)
    If $\Psi$ is a NO instance, then $V_{\Psi}$ rejects every proof $y$.
Proof.
We describe the deterministic quantum reduction that maps $\Psi$ to a verifier $V_{\Psi}$. The following data is hardcoded into the description of $V_{\Psi}$:
$\left\{S_{j}\subseteq[m]:|S_{j}|\leq d\right\},\quad\{\beta_{S_{j}}\}_{j=1}^{N},\quad a,\,\delta_{D},$
where $N=\sum_{i=0}^{d}\binom{m}{i}=\mathrm{poly}(m)$.
Verifier $V_{\Psi}$.
On input proof $y\in\{0,1\}^{m}$, the verifier proceeds as follows:
  1. (i)
    Compute $B:=\sum_{j=1}^{N}|\beta_{S_{j}}|$ and sample a uniform real $r\in[0,B)$ using $\lceil\log(n)\rceil$ random bits.
  2. (ii)
    Find the unique index $j$ such that
    $\sum_{i=1}^{j-1}|\beta_{S_{i}}|\leq r<\sum_{i=1}^{j}|\beta_{S_{i}}|.$ (2.24)
    Let $S=S_{j}$ and $\sigma:=\mathrm{sign}(\beta_{S_{j}})\in\{\pm 1\}$.
  3. (iii)
    The verifier non-adaptively queries the qubits $y_{i}$ for $i\in S$ into an ancilla register via the unitary:
    $|i\rangle|0\rangle\mapsto|i\rangle|y_{i}\rangle,\quad\forall i\in S.$ (2.25)
  4. (iv)
    Phase kickback. Apply a controlled-phase gate that multiplies the global phase by
    $\exp\left(i\pi\sigma\prod_{i\in S}y_{i}\right).$ (2.26)
  5. (v)
    Estimate contribution. Use a Hadamard test subroutine on the ancilla to estimate the real part $\Re[\beta_{S}\prod_{i\in S}y_{i}]$ up to additive error $\delta_{D}/2$.
  6. (vi)
    To ensure this accuracy with high probability, repeat the Hadamard test subroutine $T=O(1/\delta_{D}^{2})$ times. This follows from Hoeffding’s inequality: given bounded random variables $X_{1},\dots,X_{T}$ with $|X_{i}|\leq 1$ and mean $\mu$, the empirical mean $\hat{\mu}=\frac{1}{T}\sum_{i=1}^{T}X_{i}$ satisfies
    $\Pr[\,|\hat{\mu}-\mu|\geq\epsilon\,]\leq 2\exp(-2T\epsilon^{2}).$
    Setting $\epsilon=\delta_{D}/2$, it suffices to take
    $T\geq\frac{2\ln(2/\delta)}{\delta_{D}^{2}}$
    for failure probability $\delta$. Since $\delta_{D}$ is constant, this requires only a constant number of repetitions.
  7. (vii)
    Accept/reject. Let $\widehat{P}(y)$ be the empirical estimate of $P(y)$ obtained by averaging the Hadamard test results over repeated monomial sampling. The verifier accepts if and only if $\widehat{P}(y)\geq a$.
    To ensure correctness with respect to the $\mathsf{QCPCP}$ gap, we normalize the thresholds by $B=\sum_{j=1}^{N}|\beta_{S_{j}}|$, setting
    $c=\frac{a+\delta_{D}}{B},\qquad s=\frac{a-\delta_{D}}{B},$
    which satisfies $c-s={2\delta_{D}}/{B}={1}/{\mathrm{poly}(m)}$, sufficient for our $\mathsf{QCPCP}$ setting.
Analysis.
With this analysis, we conclude that $\mathsf{UniqueMTP}\leq_{\mathsf{BQ}}\mathsf{UniqueQCPCP}$.
  1. (i)
    (Bounded query) In each execution, $V_{\Psi}$ queries at most $|S|\leq d$ bits of the proof $y$, thus the query complexity is $q=d$.
  2. (ii)
    (Completeness) If $\Psi$ is a YES instance with unique satisfying assignment $y^{*}$ such that $P(y^{*})\geq a$, then each monomial contributes exactly $\beta_{S}\prod_{i\in S}y^{*}_{i}$, so $\mathbb{E}[\widehat{P}(y^{*})]=P(y^{*})\geq a$. By concentration bounds, the empirical estimate $\widehat{P}(y^{*})$ exceeds $a$ with high probability. No other $y\neq y^{*}$ can satisfy $P(y)\geq a$, hence only $y^{*}$ is accepted.
  3. (iii)
    (Soundness) If $\Psi$ is a NO instance, then for all $y\in\{0,1\}^{m}$, we have $P(y)<a$, so with high probability the estimate $\widehat{P}(y)$ also satisfies $\widehat{P}(y)<a$, and $V_{\Psi}$ rejects all proofs.
  4. (iv)
    (Uniqueness preservation) The reduction $\Psi\mapsto V_{\Psi}$ preserves the uniqueness of the satisfying assignment: the only proof that can be accepted by $V_{\Psi}$ is the unique $y$ such that $P(y)\geq a$.
This completes the proof. ∎
Theorem 2.11 (Reduction from $\mathsf{QCPCP}$ to $\mathsf{UniqueQCPCP}$).
Let $V$ be a $\mathsf{QCPCP}$ verifier for a promise problem $\mathcal{A}=(A_{\mathrm{YES}},A_{\mathrm{NO}})$, with completeness $c$, soundness $s$, proof length $l_{n}$, and running time $T_{n}$. Then there exists a randomized polynomial-time algorithm which, given $V$, outputs a verifier $V^{\prime}$ for a $\mathsf{UniqueQCPCP}$ system with the following properties:
  1. (i)
    $V^{\prime}$ runs in time $\mathcal{O}(T_{n})$ and expects proofs of length $l_{n}+\mathcal{O}(\log(l_{n}))$.
  2. (ii)
    (Completeness): If $x\in A_{\mathrm{YES}}$, then with probability at least $\Omega(1/\textnormal{poly}(l_{n}))$ over the randomness of the reduction, there exists a unique proof $y^{*}\in\{0,1\}^{l_{n}+\mathcal{O}(\log(l_{n}))}$ such that:
    $\Pr[V^{\prime}(x,y^{*})=1]\geq c-\gamma,\quad\text{and}\quad\forall y^{\prime}\neq y^{*},\;\Pr[V^{\prime}(x,y^{\prime})=1]\leq\gamma,$ (2.27)
    where $\gamma=\operatorname{negl}(n)$ is negligible in $n$.
  3. (iii)
    (Soundness) If $x\in A_{\mathrm{NO}}$, then with probability $1$ over the randomness of the reduction, it holds that for all $y\in\{0,1\}^{l_{n}+\mathcal{O}(\log(l_{n}))}$,
    $\Pr[V^{\prime}(x,y)=1]\leq s+\gamma.$ (2.28)
Proof.
This follows by composing the reductions:
$\mathsf{QCPCP}\leq_{\mathsf{BQ}}\mathsf{MTP}\leq_{p}\mathsf{SAT}\leq_{\mathsf{RP}}\mathsf{UniqueSAT}\leq_{p}\mathsf{UniqueMTP}\leq_{\mathsf{BQ}}\mathsf{UniqueQCPCP},$ (2.29)
where the first step is given by of [BGW24, Theorem 4], and the remaining steps are covered in Lemma 2.8, Valiant–Vazirani reduction [VV86]Lemma 2.9, and Lemma 2.10. The only degradation in completeness and soundness arises from the randomized reduction from $\mathsf{SAT}$ to $\mathsf{UniqueSAT}$, which incurs a loss of at most $1/l_{n}$ in each. This completes the proof. ∎
2.3 Search-to-decision reductions
For $\mathsf{QCMA}$, [INN+22] established a one-query Bernstein-Vazirani-style reduction using an oracle of the form $f(x)=x\cdot y^{*}$, where $y^{*}$ is the desired witness and $x\cdot y^{*}$ denotes the inner product. This reduction exploits the fact that $\mathsf{QCMA}$ verifiers have access to the entire classical proof string.
The situation for $\mathsf{QCPCP}$ is fundamentally different. A $\mathsf{QCPCP}$ oracle can only provide information of the form “there exists a proof $y$ such that the verifier accepts,” rather than direct access to inner products with proof bits. This limitation arises because $\mathsf{QCPCP}$ verifiers query only a constant number of bits from the proof, rather than having access to the entire proof string as in $\mathsf{QCMA}$.
Consequently, witness extraction for $\mathsf{QCPCP}$ appears to require bit-by-bit reconstruction, leading to $\mathcal{O}(n)$ oracle queries where $n$ is the proof length. With current techniques, no one-query Bernstein-Vazirani-style reduction analogous to that of [INN+22] seems feasible for $\mathsf{QCPCP}$.
3 Towards the collapse of the quantum polynomial hierarchy
In this section, we investigate two perspectives on the potential collapse of the quantum polynomial hierarchy ($\mathsf{QPH}$). First, we present a non-uniform version of the Karp–Lipton theorem for $\mathsf{QPH}$. Second, we introduce a new class of problems that is structurally closer to the standard $\mathsf{QPH}$: namely, the proofs are restricted to product states (in contrast to the entangled states allowed in $\mathsf{QEPH}$), but are still interdependent due to a consistency condition enforced by the interaction of two players.
3.1 Definitions and preliminaries
3.1.1 Complexity classes
We begin by introducing the necessary definitions for quantum complexity classes including $\mathsf{Q\Sigma}_{i}(c,s)$, $\mathsf{QPH}$, $\mathsf{BQP}/\mathsf{qpoly}$ and $\mathsf{Q\Sigma}_{2}/\mathsf{qpoly}$.
Definition 3.1 ($\mathsf{Q\Sigma}_{i}(c,s)$).
Let $\mathcal{A}=(A_{\mathrm{YES}},A_{\mathrm{NO}})$ be a promise problem, and let $c,s:\mathbb{N}\rightarrow[0,1]$ be polynomial-time computable functions. Then $\mathcal{A}\in\mathsf{Q\Sigma}_{i}(c,s)$ if there exists:
  1. (i)
    a polynomially bounded function $p:\mathbb{N}\to\mathbb{N}$, and
  2. (ii)
    a polynomial-time uniform family of quantum circuits $\{V_{n}\}_{n\in\mathbb{N}}$,
such that for every input $x\in\{0,1\}^{n}$, the verifier $V_{n}$ takes $p(n)$-qubit quantum states $\rho_{1},\ldots,\rho_{i}$ as input proofs and outputs a single qubit. The verifier measures the output qubit in the computational basis, and the acceptance probability is defined on input $\rho_{1}\otimes\cdots\otimes\rho_{i}$.
  1. (i)
    (Completeness) If $x\in L_{\mathrm{YES}}$, then
    $\exists\rho_{1}\;\forall\rho_{2}\;\cdots\;Q_{i}\rho_{i}\quad\textnormal{s.t.}\quad\Pr[V_{n}\textnormal{ accepts}]\geq c(n).$ (3.1)
  2. (ii)
    (Soundness) If $x\in L_{\mathrm{NO}}$, then
    $\forall\rho_{1}\;\exists\rho_{2}\;\cdots\;\overline{Q}_{i}\rho_{i}\quad\textnormal{s.t.}\quad\Pr[V_{n}\textnormal{ accepts}]\leq s(n).$ (3.2)
Here, $Q_{i}=\forall$ if $i$ is even and $Q_{i}=\exists$ if $i$ is odd; $\overline{Q}_{i}$ denotes the complementary quantifier. When the completeness and soundness parameters are not specified, we define the class
$\mathsf{Q\Sigma}_{i}:=\bigcup_{c-s\in\Omega(1/\mathrm{poly}(n))}\mathsf{Q\Sigma}_{i}(c,s).$ (3.3)
Definition 3.2 (Quantum polynomial hierarchy, $\mathsf{QPH}$).
The quantum polynomial hierarchy $\mathsf{QPH}$ is defined as
$\mathsf{QPH}:=\bigcup_{i\in\mathbb{N}}\mathsf{Q\Sigma}_{i}=\bigcup_{i\in\mathbb{N}}\mathsf{Q\Pi}_{i},$ (3.4)
where $\mathsf{Q\Pi}_{i}:=\mathsf{co\text{-}Q\Sigma}_{i}$.
Definition 3.3 ($\mathsf{BQP}/\mathsf{qpoly}$).
A promise problem $\mathcal{A}=(A_{\mathrm{YES}},A_{\mathrm{NO}})$ is in $\mathsf{BQP}/\mathsf{qpoly}$ if there exist:
  1. (i)
    a polynomial-size family of quantum circuits $\{C_{n}\}_{n\in\mathbb{N}}$, and
  2. (ii)
    a sequence of polynomial-size quantum advice states $\{|\psi_{n}\rangle\}_{n\in\mathbb{N}}$,
such that for all $x\in\{0,1\}^{n}$, the circuit $C_{n}$ takes input $|x\rangle\otimes|\psi_{n}\rangle$ and satisfies:
  1. (i)
    (Completeness) If $x\in A_{\mathrm{YES}}$, then
    $\Pr[C_{n}(|x\rangle\otimes|\psi_{n}\rangle)=1]\geq 2/3.$ (3.5)
  2. (ii)
    (Soundness) If $x\in A_{\mathrm{NO}}$, then
    $\Pr[C_{n}(|x\rangle\otimes|\psi_{n}\rangle)=1]\leq 1/3.$ (3.6)
Note that $\mathsf{BQP}/\mathsf{qpoly}$ is the quantum analogue of $\mathsf{BPP}/\mathsf{poly}$, where quantum advice states replace classical advice strings.
Definition 3.4 ($\mathsf{Q\Sigma}_{2}/\mathsf{qpoly}$).
A promise problem $\mathcal{A}=(A_{\mathrm{YES}},A_{\mathrm{NO}})$ is in $\mathsf{Q\Sigma}_{2}/\mathsf{qpoly}(c,s)$ if there exist:
  1. (i)
    a polynomial-time uniform family of quantum circuits $\{V_{n}\}_{n\in\mathbb{N}}$,
  2. (ii)
    a sequence of polynomial-size quantum advice states $\{|\psi_{n}\rangle\}_{n\in\mathbb{N}}$, and
  3. (iii)
    polynomial-time computable functions $c,s:\mathbb{N}\rightarrow[0,1]$,
such that for all $x\in\{0,1\}^{n}$, the verifier $V_{n}$ takes $(x,|\psi_{n}\rangle,\rho_{1},\rho_{2})$ as input, where $\rho_{1},\rho_{2}$ are polynomial-size quantum states, and satisfies:
  1. (i)
    (Completeness) If $x\in A_{\mathrm{YES}}$, then
    $\exists\rho_{1}\;\forall\rho_{2}\quad\text{s.t.}\quad\Pr[V_{n}(x,|\psi_{n}\rangle,\rho_{1},\rho_{2})=1]\geq c(n).$ (3.7)
  2. (ii)
    (Soundness) If $x\in A_{\mathrm{NO}}$, then
    $\forall\rho_{1}\;\exists\rho_{2}\quad\text{s.t.}\quad\Pr[V_{n}(x,|\psi_{n}\rangle,\rho_{1},\rho_{2})=1]\leq s(n).$ (3.8)
When the completeness-soundness gap is not explicitly given, we define
$\mathsf{Q\Sigma}_{2}/\mathsf{qpoly}:=\bigcup_{c-s\in\Omega(1/\mathrm{poly}(n))}\mathsf{Q\Sigma}_{2}/\mathsf{qpoly}(c,s).$ (3.9)
Definition 3.5 (Quantum entangled polynomial hierarchy, $\mathsf{QEPH}$ [GY24]).
The class $\mathsf{QEPH}$ is defined analogously to $\mathsf{QPH}$, except that the proof states $\rho_{1},\ldots,\rho_{i}$ in the completeness and soundness conditions are allowed to be arbitrary entangled states over all proof registers, rather than being restricted to product states as in $\mathsf{QPH}$.
Remark 3.6 (Containment relations).
It is straightforward to verify that $\mathsf{QPH}\subseteq\mathsf{QEPH}$ and $\mathsf{BQP}/\mathsf{qpoly}\subseteq\mathsf{Q\Sigma}_{2}/\mathsf{qpoly}$. The relationship between quantum and classical advice in the hierarchy setting (e.g., whether $\mathsf{Q\Sigma}_{i}/\mathsf{qpoly}\subseteq\mathsf{Q\Sigma}_{i}/\mathsf{poly}$) remains an interesting open question.
3.1.2 Convexity considerations in quantum proof systems
In this section, we describe the convexity-related properties that are essential for proving our collapse result.
Theorem 3.7 (Sion’s minimax [Sio58]).
Let $X$ and $Y$ be topological vector spaces, and let $A\subseteq X$ and $B\subseteq Y$ be convex and compact subsets. Suppose $f:A\times B\rightarrow\mathbb{R}$ is a function that is concave in its first argument and convex in its second argument. Then,
$\max_{a\in A}\min_{b\in B}f(a,b)=\min_{b\in B}\max_{a\in A}f(a,b).$ (3.10)
In the context of quantum proof systems, the set of density matrices $\mathcal{D}(\mathcal{H})$ over a finite-dimensional Hilbert space $\mathcal{H}$ is convex and compact in the trace norm topology [Wat18].
Theorem 3.8 (Krein–Milman [MK40]).
Let $K$ be a non-empty compact convex subset of a locally convex topological vector space $X$. Then
$K=\overline{\mathrm{conv}}(\operatorname{ext}(K)),$
where $\operatorname{ext}(K)$ denotes the set of extreme points of $K$, and $\overline{\mathrm{conv}}$ denotes the closure of the convex hull with respect to the topology of $X$.
Lemma 3.9 (Convexity of consistent mixed product states and non-convexity of consistent pure product states).
Let $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ be finite-dimensional Hilbert spaces. Then:
  1. (i)
    The set of consistent (mixed) product states
    $\mathcal{S}(\rho_{1}):=\left\{\rho\in\mathcal{D}(\mathcal{H}_{1}\otimes\mathcal{H}_{2}):\rho=\rho_{1}\otimes\rho_{2}\textnormal{ for some }\rho_{2}\in\mathcal{D}(\mathcal{H}_{2})\right\}$ (3.11)
    is convex for any fixed $\rho_{1}\in\mathcal{D}(\mathcal{H}_{1})$.
  2. (ii)
    The set of consistent pure product states
    $\mathcal{S}_{\mathrm{pure}}(\psi_{1}):=\left\{\rho=|\psi_{1}\rangle\langle\psi_{1}|\otimes|\psi_{2}\rangle\langle\psi_{2}|:|\psi_{2}\rangle\in\mathcal{H}_{2}\right\}$ (3.12)
    is not convex for any fixed pure state $|\psi_{1}\rangle\in\mathcal{H}_{1}$.
Proof.
Let $\rho_{1}\in\mathcal{D}(\mathcal{H}_{1})$, and consider the set $\mathcal{S}(\rho_{1})=\left\{\rho=\rho_{1}\otimes\rho_{2}:\rho_{2}\in\mathcal{D}(\mathcal{H}_{2})\right\}$.
Let $\rho^{(i)}=\rho_{1}\otimes\rho_{2}^{(i)}\in\mathcal{S}(\rho_{1})$ for $i=1,\dots,m$, and let $\lambda_{1},\dots,\lambda_{m}$ be non-negative weights summing to 1. Then their convex combination is
$\rho^{\prime}=\sum_{i=1}^{m}\lambda_{i}\rho^{(i)}=\sum_{i=1}^{m}\lambda_{i}\left(\rho_{1}\otimes\rho_{2}^{(i)}\right)=\rho_{1}\otimes\left(\sum_{i=1}^{m}\lambda_{i}\rho_{2}^{(i)}\right).$ (3.13)
Since the convex combination of density operators is again a density operator, it follows that $\rho^{\prime}\in\mathcal{S}(\rho_{1})$. Thus, $\mathcal{S}(\rho_{1})$ is convex.
Now consider a fixed pure state $|\psi_{1}\rangle\in\mathcal{H}_{1}$, and take $\rho^{(i)}=|\psi_{1}\rangle\langle\psi_{1}|\otimes|\psi_{2}^{(i)}\rangle\langle\psi_{2}^{(i)}|\in\mathcal{S}_{\mathrm{pure}}(\psi_{1})$ for $i=1,\dots,m$, and form the convex combination:
$\rho^{\prime}=\sum_{i=1}^{m}\lambda_{i}\rho^{(i)}=|\psi_{1}\rangle\langle\psi_{1}|\otimes\left(\sum_{i=1}^{m}\lambda_{i}|\psi_{2}^{(i)}\rangle\langle\psi_{2}^{(i)}|\right),$ (3.14)
for some convex weights $\lambda_{i}\geq 0$, $\sum_{i}\lambda_{i}=1$. The second subsystem is now in a mixed state unless all $|\psi_{2}^{(i)}\rangle$ are equal. Hence, $\rho^{\prime}\notin\mathcal{S}_{\mathrm{pure}}(\psi_{1})$ in general, and so the set is not convex. ∎
3.2 Non-uniform Karp–Lipton collapse for quantum polynomial hierarchy
In this section, we prove the following quantum Karp–Lipton theorem by leveraging the quantum complexity classes defined in the previous section along with key convexity properties.
Theorem 3.10 (Quantum Karp–Lipton).
If $\mathsf{QMA}\subseteq\mathsf{BQP}/\mathsf{qpoly}$, then the quantum polynomial hierarchy collapses to its second level with quantum advice:
$\mathsf{QPH}=\bigcup_{k\in\mathbb{N}}\mathsf{Q\Sigma}_{k}\subseteq\mathsf{Q\Sigma}_{2}/\mathsf{qpoly}.$ (3.15)
Proof.
We prove the result by induction on $k$. Throughout the proof, we use the standard completeness and soundness parameters $c=2/3$ and $s=1/3$.
Base Case ($k=3$).
Let $\mathcal{A}\in\mathsf{Q\Sigma_{3}}$. Then by Definition 3.1, there exists a polynomial-time uniform family of quantum circuits $\{V_{n}^{(3)}\}$ such that for all $x\in\{0,1\}^{n}$:
$x\in A_{\mathrm{YES}}$ $\iff\exists\rho_{1}\,\forall\rho_{2}\,\exists\rho_{3}:\Pr[V_{n}^{(3)}(x;\rho_{1},\rho_{2},\rho_{3})=1]\geq 2/3,$ (3.16)
$x\in A_{\mathrm{NO}}$ $\iff\forall\rho_{1}\,\exists\rho_{2}\,\forall\rho_{3}:\Pr[V_{n}^{(3)}(x;\rho_{1},\rho_{2},\rho_{3})=1]\leq 1/3.$ (3.17)
Define a quantum promise problem $\mathsf{P}_{x}$ with input $(\rho_{1},\rho_{2})$ as follows:
$(\rho_{1},\rho_{2})\in\mathsf{P}_{x,\text{YES}}$ $\iff\exists\rho_{3}:\Pr[V_{n}^{(3)}(x;\rho_{1},\rho_{2},\rho_{3})=1]\geq 2/3,$ (3.18)
$(\rho_{1},\rho_{2})\in\mathsf{P}_{x,\text{NO}}$ $\iff\forall\rho_{3}:\Pr[V_{n}^{(3)}(x;\rho_{1},\rho_{2},\rho_{3})=1]\leq 1/3.$ (3.19)
By construction, $\mathsf{P}_{x}\in\mathsf{QMA}$ for each fixed $x$. By the assumption $\mathsf{QMA}\subseteq\mathsf{BQP}/\mathsf{qpoly}$, there exists a polynomial-time quantum circuit family $\{C_{n}\}$ and a polynomial-size quantum advice sequence $\{|\alpha_{n}\rangle\}$ (depending only on $n=|x|$) such that:
$(\rho_{1},\rho_{2})\in\mathsf{P}_{x,\text{YES}}$ $\implies\Pr[C_{n}(x,\rho_{1},\rho_{2},|\alpha_{n}\rangle)=1]\geq 2/3,$ (3.20)
$(\rho_{1},\rho_{2})\in\mathsf{P}_{x,\text{NO}}$ $\implies\Pr[C_{n}(x,\rho_{1},\rho_{2},|\alpha_{n}\rangle)=1]\leq 1/3.$ (3.21)
Substituting this characterization into the original quantifier expressions:
$x\in A_{\mathrm{YES}}$ $\iff\exists\rho_{1}\,\forall\rho_{2}:\Pr[C_{n}(x,\rho_{1},\rho_{2},|\alpha_{n}\rangle)=1]\geq 2/3,$ (3.22)
$x\in A_{\mathrm{NO}}$ $\iff\forall\rho_{1}\,\exists\rho_{2}:\Pr[C_{n}(x,\rho_{1},\rho_{2},|\alpha_{n}\rangle)=1]\leq 1/3.$ (3.23)
This defines a $\mathsf{Q\Sigma_{2}}$ protocol with quantum advice $|\alpha_{n}\rangle$. Therefore, $\mathcal{A}\in\mathsf{Q\Sigma_{2}}/\mathsf{qpoly}$.
Inductive Step.
Assume the conditional inclusion holds for all $\mathsf{Q\Sigma}_{j}$ with $j\leq k-1$, i.e., $\mathsf{Q\Sigma}_{j}\subseteq\mathsf{Q\Sigma_{2}}/\mathsf{qpoly}$. We prove that $\mathsf{Q\Sigma}_{k}\subseteq\mathsf{Q\Sigma_{2}}/\mathsf{qpoly}$ for $k\geq 4$.
Let $\mathcal{A}\in\mathsf{Q\Sigma}_{k}$. By definition:
$x\in A_{\mathrm{YES}}\iff\exists\rho_{1}\,\forall\rho_{2}\,\exists\rho_{3}\cdots Q_{k}\rho_{k}:\Pr[V_{n}^{(k)}(x;\rho_{1},\ldots,\rho_{k})=1]\geq 2/3.$ (3.24)
We can restructure this as:
$x\in A_{\mathrm{YES}}\iff\exists\rho_{1}\,\forall\rho_{2}:(x,\rho_{1},\rho_{2})\in\mathcal{A}^{\prime},$ (3.25)
where the quantum language $\mathcal{A}^{\prime}\in\mathsf{Q\Sigma}_{k-2}$ is defined by:
$(x,\rho_{1},\rho_{2})\in A^{\prime}_{\mathrm{YES}}\iff\exists\rho_{3}\cdots Q_{k}\rho_{k}:\Pr[V_{n}^{(k)}(x;\rho_{1},\ldots,\rho_{k})=1]\geq 2/3.$ (3.26)
By the induction hypothesis, $\mathcal{A}^{\prime}\in\mathsf{Q\Sigma_{2}}/\mathsf{qpoly}$. Therefore, there exists a verifier $N_{n}$ and advice state $|\beta_{n}\rangle$ such that:
$(x,\rho_{1},\rho_{2})\in A^{\prime}_{\mathrm{YES}}\iff\exists\sigma_{1}\,\forall\sigma_{2}:\Pr[N_{n}(x,\rho_{1},\rho_{2},\sigma_{1},\sigma_{2},|\beta_{n}\rangle)=1]\geq 2/3.$ (3.27)
Substituting back:
$x\in A_{\mathrm{YES}}\iff\exists\rho_{1}\,\forall\rho_{2}\,\exists\sigma_{1}\,\forall\sigma_{2}:\Pr[N_{n}(x,\rho_{1},\rho_{2},\sigma_{1},\sigma_{2},|\beta_{n}\rangle)=1]\geq 2/3.$ (3.28)
The inner three quantifiers $\forall\rho_{2}\exists\sigma_{1}\forall\sigma_{2}$ form a $\mathsf{Q\Pi_{3}/qpoly}$ predicate. Using base case and $\mathsf{Q\Sigma_{2}=Q\Pi_{2}}$, we have
$\mathsf{Q\Pi}_{3}/\mathsf{qpoly}$ $=\mathsf{co{-}(Q\Sigma}_{3}/\mathsf{qpoly)}\subseteq\mathsf{co{-}(Q\Sigma}_{2}/\mathsf{qpoly)}$ (3.29)
$=\mathsf{Q\Pi}_{2}/\mathsf{qpoly}$ (3.30)
$=\mathsf{Q\Sigma}_{2}/\mathsf{qpoly}.$ (3.31)
Thus this predicate can be squeezed to a $\mathsf{Q\Sigma_{2}/qpoly}$ as follows, where $M_{n}$ is a $\mathsf{Q\Sigma_{2}/qpoly}$ verifier and quantum promise language $\mathcal{A}^{\prime\prime}\in\mathsf{Q\Pi_{3}/qpoly}$.
$(x,\rho_{1})\in A_{\textnormal{YES}}^{\prime\prime}\iff\forall\rho_{2}\exists\sigma_{1}\forall\sigma_{2}:\operatorname{Pr}[N_{n}(x,\rho_{1},\rho_{2},\sigma_{1},\sigma_{2},|\beta_{n}\rangle)=1]\geq 2/3.$ (3.32)
$(x,\rho_{1})\in A^{\prime\prime}_{\textnormal{YES}}\iff\exists\rho_{1}^{\prime}\forall\sigma_{1}^{\prime}:\operatorname{Pr}[M_{n}(x,\rho_{1},\rho_{1}^{\prime},\sigma_{1}^{\prime},|\gamma_{n}\rangle)=1]\geq 2/3.$ (3.33)
Substituting back,
$x\in A_{\textnormal{YES}}\iff\exists\rho_{1}\exists\rho_{1}^{\prime}\forall\sigma_{1}^{\prime}:\operatorname{Pr}[M_{n}(x,\rho_{1},\rho_{1}^{\prime},\sigma_{1}^{\prime},|\gamma_{n}\rangle)=1]\geq 2/3,$ (3.34)
which is a $\mathsf{Q\Sigma_{2}/qpoly}$ predicate with existential proof $\rho_{1}\otimes\rho_{1}^{\prime}$ and universal proof $\sigma_{1}^{\prime}$. The soundness condition follows by similar reasoning.
Hence, $\mathsf{Q\Sigma}_{k}\subseteq\mathsf{Q\Sigma_{2}}/\mathsf{qpoly}$ for all $k\geq 3$, and the result follows for $k\leq 2$ trivially. ∎
3.3 Consistent quantum polynomial hierarchy
We introduce a variant of the quantum polynomial hierarchy that enforces structural consistency constraints across rounds of interaction. This variant, which we denote by $\mathsf{CQPH}$, differs from the standard $\mathsf{QPH}$ in that it requires provers to maintain coherent extensions of their quantum states throughout the protocol execution.
Definition 3.11 (Consistent extension property).
Let $\mathcal{H}_{1},\mathcal{H}_{2},\ldots,\mathcal{H}_{k}$ be finite-dimensional Hilbert spaces. A sequence of quantum states $\{\rho_{i}\}_{i=1}^{k}$ satisfies the consistent extension property if for each $i\geq 2$, there exists a state $\rho_{i}$ on $\mathcal{H}_{1}\otimes\cdots\otimes\mathcal{H}_{i}$ such that $\mathrm{Tr}_{\mathcal{H}_{i}}(\rho_{i})=\rho_{i-1}$.
In a $\mathsf{CQPH}$ protocol, each prover must submit states that satisfy this property. Specifically, if Alice’s $i$-th message is a state on registers $\mathsf{A}_{1}\otimes\cdots\otimes\mathsf{A}_{i}$, then tracing out $\mathsf{A}_{i}$ must yield her $(i{-}1)$-th message exactly.
Definition 3.12 (Consistent quantum polynomial hierarchy, $\mathsf{CQPH}$).
Let $c,s:\mathbb{N}\to[0,1]$ be polynomial-time computable functions with $c(n)-s(n)\geq 1/\mathrm{poly}(n)$, where $n=|x|$. A promise problem $\mathcal{A}=(A_{\mathrm{YES}},A_{\mathrm{NO}})$ belongs to $\mathsf{CQ\Sigma}_{k}(c,s)$ if there exists a polynomial-time uniform family of quantum circuits $\{V_{n}\}$ such that:
  1. (i)
    (Completeness) For $x\in A_{\mathrm{YES}}$, there exist consistent quantum proofs achieving
    $\exists\rho_{1}\,\forall\sigma_{1}\,\exists\rho_{2}\,\forall\sigma_{2}\,\cdots\,Q_{k}\tau_{k}:\Pr[V_{n}(\rho_{1},\sigma_{1},\ldots,\tau_{k})=1]\geq c(n)$ (3.35)
  2. (ii)
    (Soundness) For $x\in A_{\mathrm{NO}}$, all consistent proof strategies satisfy
    $\forall\rho_{1}\,\exists\sigma_{1}\,\forall\rho_{2}\,\exists\sigma_{2}\,\cdots\,\overline{Q}_{k}\tau_{k}:\Pr[V_{n}(\rho_{1},\sigma_{1},\ldots,\tau_{k})=1]\leq s(n)$ (3.36)
where $Q_{k}=\exists$ if $k$ is odd and $Q_{k}=\forall$ if $k$ is even, and all states must satisfy the consistent extension property within their respective proof sequences. We define $\mathsf{CQ\Sigma}_{k}=\bigcup_{c,s}\mathsf{CQ\Sigma}_{k}(c,s)$ and $\mathsf{CQPH}=\bigcup_{k=1}^{\infty}\mathsf{CQ\Sigma}_{k}$.
Definition 3.13 (Pure-state variant of $\mathsf{CQPH}$ and $\mathsf{CQ\Sigma}_{i}$).
We define the complexity class $\mathsf{pure\text{-}CQPH}$ analogously to $\mathsf{CQPH}$, with the distinction that all quantum proofs are required to be pure states over their respective Hilbert spaces. That is, the verifier must accept a collection of pure quantum states as input proofs, rather than general (possibly mixed) quantum states. Similarly, we define $\mathsf{pure\text{-}CQ\Sigma}_{i}$ as the variant of $\mathsf{CQ\Sigma}_{i}$ in which all quantum proofs are restricted to be pure states.
The collapse results we derive for the newly defined consistent variant complexity classes are formally stated as follows.
Theorem 3.14 (Collapse of $\mathsf{CQPH}$ and $\mathsf{pure\text{-}CQPH}$).
$\mathsf{CQPH}=\mathsf{CQ\Sigma}_{2}$ and $\mathsf{pure\text{-}CQPH}=\mathsf{pure\text{-}CQ\Sigma}_{2}$.
Proof.
We establish that $\mathsf{CQ\Sigma}_{3}\subseteq\mathsf{CQ\Sigma}_{2}$ (respectively, $\mathsf{pure\text{-}CQ\Sigma}_{3}\subseteq\mathsf{pure\text{-}CQ\Sigma}_{2}$), and extend the result inductively to all higher levels.
Consider a $\mathsf{CQ\Sigma}_{3}$ protocol whose acceptance probability is characterized by the following min-max expression:
$v^{*}=\max_{\rho_{1}\in\mathcal{D}(\mathsf{A}_{1})}\min_{\sigma_{1}\in\mathcal{D}(\mathsf{B}_{1})}\max_{\rho_{2}\in\mathcal{S}(\rho_{1})}\operatorname{Tr}(R(\rho_{2}\otimes\sigma_{1})),$ (3.37)
where $\mathcal{S}(\rho_{1}):=\left\{\rho=\rho_{1}\otimes\rho^{\prime}\in\mathcal{D}(\mathsf{A}_{1}\otimes\mathsf{A}_{2}):\operatorname{Tr}_{\mathsf{A}_{2}}(\rho)=\rho_{1},~{}\rho^{\prime}\in\mathcal{D}(\mathsf{A}_{2})\right\}$ is the set of consistent extensions of $\rho_{1}$.
In the pure-state setting, the analogous quantity is given by:
$v^{*}_{\text{pure}}=\max_{|\psi_{1}\rangle\in\mathsf{A}_{1}}\min_{|\phi_{1}\rangle\in\mathsf{B}_{1}}\max_{\rho_{2}\in\mathcal{S}_{\text{pure}}(|\psi_{1}\rangle)}\operatorname{Tr}(R(\rho_{2}\otimes\sigma_{1})),$ (3.38)
where $\mathcal{S}_{\text{pure}}(|\psi_{1}\rangle):=\left\{\rho=|\psi_{1}\rangle\langle\psi_{1}|\otimes|\psi^{\prime}\rangle\langle\psi^{\prime}|\in\mathcal{D}(\mathsf{A_{1}}\otimes\mathsf{A_{2}}):\operatorname{Tr}_{\mathsf{A}_{2}}(\rho)=|\psi_{1}\rangle\langle\psi_{1}|\right\}$.
As established in Lemma 3.9, $\mathcal{S}(\rho_{1})$ forms a convex set, whereas $\mathcal{S}_{\mathrm{pure}}(|\psi_{1}\rangle)$ does not. However, each element of $\mathcal{S}_{\text{pure}}(|\psi_{1}\rangle)$ is an extremal point of its convex hull. By the Krein–Milman theorem (Theorem 3.8), it suffices to optimize over $\operatorname{conv}(\mathcal{S}_{\text{pure}}(|\psi_{1}\rangle))$.
Hence, we may equivalently write:
$v^{*}=\max_{\rho_{1}\in\mathcal{D}(\mathsf{A_{1}})}\max_{\rho_{2}\in\mathcal{S}(\rho_{1})}\min_{\sigma_{1}\in\mathcal{D}(\mathsf{B_{1}})}\operatorname{Tr}(R(\rho_{2}\otimes\sigma_{1})),$ (3.39)
and
$v^{*}_{\text{pure}}=\max_{|\psi_{1}\rangle}\max_{\rho_{2}\in\operatorname{conv}(\mathcal{S}_{\text{pure}}(|\psi_{1}\rangle))}\min_{|\phi_{1}\rangle}\operatorname{Tr}(R(\rho_{2}\otimes\sigma_{1})).$ (3.40)
In both cases, the innermost optimization is over convex sets, and the objective is bilinear in the optimization variables. Thus, we may invoke Sion’s minimax theorem to exchange the min and max operators, reducing the overall expression to a two-alternation form, i.e., a $\mathsf{CQ\Sigma}_{2}$ (resp. $\mathsf{pure\text{-}CQ\Sigma}_{2}$) protocol.
The inductive step to collapse higher levels of the hierarchy follows analogously by applying the same convexity and extremal point arguments to each successive quantifier block. ∎
Corollary 3.15 (Equivalence of quantum interactive classes).
$\mathsf{QEPH}=\mathsf{QE\Sigma}_{2}=\mathsf{CQPH}=\mathsf{CQ\Sigma}_{2}=\mathsf{QRG}(1).$ (3.41)
Proof.
The equality $\mathsf{QEPH}=\mathsf{QE\Sigma}_{2}$ follows from [GY24]. By Theorem 3.14, we have $\mathsf{CQPH}=\mathsf{CQ\Sigma}_{2}$. The equivalence with $\mathsf{QRG}(1)$ follows from the known correspondence between constant-round quantum refereed games and second-level quantum interactive proof systems [JW09]. ∎
3.4 Structural comparison of hierarchy collapses
Corollary 3.15 establishes that both $\mathsf{QEPH}$ and $\mathsf{CQPH}$ collapse to their second levels, despite having fundamentally different structural properties. This provides new insight into the mechanisms underlying hierarchy collapse in quantum interactive proof systems.
In $\mathsf{QEPH}$, the collapse mechanism relies crucially on the ability of provers to share entanglement across rounds [GY24]. The consistency requirements in that setting allow for quantum correlations that effectively “short-circuit” higher-level quantifier alternations.
In contrast, $\mathsf{CQPH}$ prohibits such entanglement sharing by construction—each prover’s state must be a consistent extension of their previous messages through local operations. Nevertheless, Theorem 3.14 shows that collapse occurs through a purely classical mechanism: the convexity of the feasible region combined with linearity of the objective function.
This dichotomy suggests that hierarchy collapse in quantum complexity theory may be more robust than previously understood, occurring through multiple independent mechanisms rather than relying solely on quantum entanglement effects. The fact that both entangled and non-entangled variants of quantum polynomial hierarchies collapse to the same level indicates a deeper structural principle at work, which merits further investigation.
Implications for quantum complexity theory.
The collapse of $\mathsf{CQPH}$ through classical optimization techniques, as opposed to quantum entanglement, suggests that certain aspects of quantum complexity may be more amenable to classical analysis than previously thought. This observation could inform the design of new quantum protocols and provide insight into the fundamental limits of quantum advantage in interactive settings.
4 Open problems and future directions
Our results naturally lead to several intriguing open questions that we believe merit further investigation. We organize these into two main categories corresponding to our primary contributions.
4.1 Uniqueness in quantum-classical probabilistically checkable proofs
The equivalence $\mathsf{UniqueQCPCP}=\mathsf{QCPCP}$ established in Section 2 opens several natural avenues for future research, particularly concerning the robustness of this collapse and its extensions to broader quantum proof systems.
Relation between $\mathsf{QPCP}$ and $\mathsf{UniqueQPCP}$.
A natural question arising from our work is whether similar results hold in the quantum proof setting as well. Currently, the only known result is a quantum oracle separation between $\mathsf{UniqueQMA}$ and $\mathsf{QMA}$, shown in [AHHN24], which itself appears to be a highly nontrivial and challenging problem.
Stronger reductions for the $\mathsf{UniqueQCPCP}=\mathsf{QCPCP}$ equivalence.
A natural direction is to investigate whether this equivalence remains true under stronger or more restricted reduction models. Our proof relies on $\mathsf{BQ}$-operator and randomized reductions, but it would be interesting to establish the equivalence under $\mathsf{BQP}$-truth-table reductions (i.e., non-adaptive quantum polynomial-time reductions), or even deterministic polynomial-time reductions. Establishing this would demonstrate that the collapse to $\mathsf{UniqueQCPCP}$ holds under more conservative assumptions about how the reduction accesses the verifier. Conversely, proving a separation would highlight fundamental structural distinctions between the two systems that are obscured by more powerful reduction models.
Extension to multi-prover and entangled variants.
Another compelling direction is to investigate whether the equivalence between $\mathsf{UniqueQCPCP}$ and $\mathsf{QCPCP}$ extends to broader classes of quantum proof systems. For instance, in the multi-prover setting, does a similar collapse hold between uniqueness and generality for $\mathsf{MIP}^{*}$-style variants of $\mathsf{QCPCP}$, where the verifier interacts with multiple (possibly entangled) provers? In such settings, uniqueness constraints may behave quite differently due to entanglement and non-local correlations, making it unclear whether the collapse still holds. Moreover, the notion of ”uniqueness” itself becomes more subtle when multiple provers are involved. Exploring these extensions could yield new insights into the role of entanglement and prover non-locality in quantum PCP constructions.
Parallel repetition for $\mathsf{UniqueQCPCP}$.
In classical PCP theory, parallel repetition theorems are crucial for amplifying soundness while preserving completeness [Raz98]. A natural question is whether there exists a quantum analogue of the parallel repetition theorem for $\mathsf{UniqueQCPCP}$ systems. Specifically, if a quantum verifier checks multiple independent copies of a $\mathsf{UniqueQCPCP}$ proof in parallel, does the soundness error decrease exponentially with the number of repetitions while completeness remains unaffected? Such a result would significantly amplify the reliability of $\mathsf{UniqueQCPCP}$ verifiers and could potentially be used to derive stronger quantum hardness-of-approximation results. However, the quantum setting presents unique challenges: entangled cheating strategies across multiple proof copies and the delicate structure of quantum measurements may disturb classical intuitions, making the existence of such a theorem highly non-trivial and potentially impactful.
4.2 Structure of the quantum polynomial hierarchy
Our investigation of $\mathsf{QPH}$ and its consistent variant $\mathsf{CQPH}$ in Section 3 reveals a rich structure that invites further exploration. Several fundamental questions remain open regarding the relationships between these hierarchies and other quantum complexity classes.
Towards a uniform Karp–Lipton theorem for $\mathsf{QPH}$.
One of the most significant open questions in quantum complexity theory is whether there exists a uniform (advice-free) Karp–Lipton collapse theorem for $\mathsf{QPH}$. Our non-uniform result shows that $\mathsf{QMA}\subseteq\mathsf{BQP}/\mathsf{qpoly}$ implies $\mathsf{QPH}\subseteq\mathsf{Q\Sigma}_{2}/\mathsf{qpoly}$, but the uniform case remains elusive. A uniform collapse would require showing that $\mathsf{QMA}\subseteq\mathsf{BQP}$ implies $\mathsf{QPH}\subseteq\mathsf{Q\Sigma}_{2}$. Such a result currently seems out of reach, as it would likely imply $\mathsf{QMA}(2)\subseteq\mathsf{PSPACE}$—a containment that is believed to be false and appears beyond the scope of existing techniques [GSS+22].
Relating $\mathsf{QPH}$ to other quantum complexity classes.
A fundamental research direction is to better understand how various levels of $\mathsf{QPH}$ relate to other well-studied quantum complexity classes such as $\mathsf{QRG}(2)$, $\mathsf{PSPACE}$, $\mathsf{QMA}(2)$, $\mathsf{QMA}$, and $\mathsf{QCMA}$. Of particular interest is whether any level of $\mathsf{QPH}$ captures the full computational power of $\mathsf{PSPACE}$. While we know that $\mathsf{QPH}\subseteq\mathsf{PSPACE}$ by definition, the reverse inclusion remains a tantalizing open problem that would provide a complete characterization of quantum hierarchical proof systems.
Formal containment of $\mathsf{CQPH}$ in $\mathsf{QPH}$.
A crucial open question is whether our consistent variant $\mathsf{CQPH}$ is strictly contained within $\mathsf{QPH}$. We strongly conjecture that $\mathsf{CQPH}\subsetneq\mathsf{QPH}$, as the consistency constraints should intuitively limit computational power. However, proving this inclusion formally remains challenging. Notably, the converse inclusion $\mathsf{QPH}\subseteq\mathsf{CQPH}$ would have dramatic consequences: since we proved $\mathsf{CQPH}=\mathsf{CQ\Sigma}_{2}\subseteq\mathsf{QRG}(1)\subseteq\mathsf{PSPACE}$ [GSS+22], this would imply an unconditional collapse of $\mathsf{QPH}$ to its second level—an outcome most researchers consider highly unlikely. Therefore, a formal proof of $\mathsf{CQPH}\subseteq\mathsf{QPH}$ would serve as important evidence against such a collapse and would have implications for the containment $\mathsf{QMA}(2)\subseteq\mathsf{PSPACE}$.
Connections between $\mathsf{CQPH}$ and $\mathsf{QRG}(2)$.
Further exploration of the relationship between $\mathsf{CQPH}$ and $\mathsf{QRG}(2)$ could illuminate the computational gap between $\mathsf{QRG}(1)$ and $\mathsf{QRG}(2)=\mathsf{PSPACE}$. Since our work establishes $\mathsf{CQPH}=\mathsf{CQ\Sigma}_{2}\subseteq\mathsf{QRG}(1)$, determining whether this inclusion is strict would clarify the landscape of quantum referee games and their relationship to hierarchical quantum proof systems. This investigation might also shed light on the power of consistency constraints in quantum interactive protocols more generally.
Comparison with oracular quantum hierarchies.
An intriguing direction is to compare $\mathsf{CQPH}$ with oracular quantum polynomial hierarchies, such as $\mathsf{QMAH}:=\mathsf{QMA}^{\mathsf{QMA}^{\mathsf{QMA}^{\cdots}}}$ (the class of problems solvable by a $\mathsf{QMA}$ machine with oracle access to $\mathsf{QMA}$, iterated polynomially many times). It remains open whether the consistency constraints in $\mathsf{CQPH}$ render it as computationally powerful as $\mathsf{QMAH}$, or whether repeated $\mathsf{QMA}$ oracle access genuinely increases computational capability. As noted in [GY24], resolving this question would provide deeper insights into the structure of $\mathsf{QRG}(1)$. Since our work establishes $\mathsf{CQPH}\subseteq\mathsf{QRG}(1)$ and it is known that $\mathsf{QMAH}\subseteq\mathsf{CH}$ [Vin18], proving $\mathsf{CQPH}\subseteq\mathsf{QMAH}$ would yield the interesting consequence $\mathsf{QRG}(1)\subseteq\mathsf{CH}$, further constraining the power of single-round quantum referee games.
Acknowledgments
J.L. acknowledges useful discussions with Nhat A. Nghiem. This work was supported by the National Research Foundation of Korea (NRF) through a grant funded by the Ministry of Science and ICT (Grant No. RS-2025-00515537). This work was also supported by the Institute for Information & Communications Technology Promotion (IITP) grant funded by the Korean government (MSIP) (Grant Nos. RS-2019-II190003 and RS-2025-02304540), the National Research Council of Science & Technology (NST) (Grant No. GTL25011-000), and the Korea Institute of Science and Technology Information (KISTI) (Grant No. P25026).